OpenStudy (anonymous):

why cos^-1(cos7pi/6)=5pi/6

OpenStudy (anonymous):

it should be 7pie/6

OpenStudy (anonymous):

its given 5pi/6 in the textbook

hartnn (hartnn):

well,cos(2pi-theta)= cos(theta) isn't it?

OpenStudy (anonymous):

cos^-1 (cos 2pie - 5pie/6) = 5pie/6

hartnn (hartnn):

careful with brackets @Yahoo!

OpenStudy (anonymous):

because \(\cos^{-1}(x)\) does not mean "any number whose cosine is \(x\)" but rather it means " the number between 0 and \(\pi\) whose cosine is \(x\)"

OpenStudy (anonymous):

since \(\frac{7\pi}{6}\) is evidently larger than \(\pi\) , it is not in the range of \(\cos^{-1}(x)\) you need another number in the interval \([0,\pi]\) with the same cosine

hartnn (hartnn):

@drown ,u still have doubts??